# Moduli stacks of Higgs bundles on stable curves

### v2

@@ 3.3.1 Relative logarithmic cotangent complex
This implies that the image of $[\sum^{n+1}_{i=1}\lambda_{i}\partial_{t_{i}}]\in T_{(1,\cdots,1)}(\mathbb{G}_{m})^{n}=\mathbb{C}^{n}$ (notice that $\sum^{n+1}_{i=1}\lambda_{i}=0$) is $\sum^{n+1}_{i=1}\lambda_{i}\cdot x_{i}\cdot\partial_{x_{i}}=\sum^{n+1}_{i=1}\lambda_{i}\cdot\partial_{log\leavevmode\nobreak\ x_{i}}$. Therefore, we get a map
$d(Or)_{p}:T_{(1,\cdots,1)}(\mathbb{G}_{m})^{n}=\mathbb{C}^{n}\longrightarrow(T\mathbb{A}^{n+1}(-log\leavevmode\nobreak\ \partial A^{n+1}))_{p}=\mathbb{C}^{n+1}$ (3.24)
$(\sum^{n+1}_{i=1}\lambda_{i}\cdot\partial_{t_{i}})\mapsto\sum^{n+1}_{i=1}\lambda_{i}\cdot\partial_{log\leavevmode\nobreak\ x_{i}}$ (3.25)
The above map is clearly injective. Therefore the dual of this map
$(d(Or)_{p})^{\vee}:\Omega^{1}_{p}\mathbb{A}^{n+1}(log\leavevmode\nobreak\ \partial A^{n+1})\longrightarrow\Omega^{1}_{(1,\cdots,1)}(\mathbb{G}_{m})^{n}$ (3.26)
@@
$\Omega_{S}^{1}(\log o)\hookrightarrow\Omega_{S[n]}^{1}(\log\leavevmode\nobreak\ \leavevmode\nobreak\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}$ (3.27)
is 00, and that the morphism $\Omega_{S[n]/S}^{1}(\log\leavevmode\nobreak\ \leavevmode\nobreak\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}$ is an isomorphism.
@@
$\big{[}\Omega_{S[n]/S}^{1}(\log\leavevmode\nobreak\ \leavevmode\nobreak\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}\big{]}$ (3.28)
is precisely the relative log cotangent complex of the morphism $\big{[}S[n]/G[n]\big{]}\longrightarrow S$ pulled back to $S[n]$. But since the morphism $\Omega_{S[n]/S}^{1}(\log\leavevmode\nobreak\ \leavevmode\nobreak\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}$ is an isomorphism, the relative log-cotangent complex of the morphism $\big{[}S[n]/G[n]\big{]}\longrightarrow S$ is equivalent to 00. ∎
3.4 Relative shifted log-symplectic forms
In this subsection, we will recall the definition of relative shifted symplectic forms for a l.f.p. morphism of derived Artin stacks. We will then define relative shifted log-symplectic forms for certain logarithmic morphisms of derived Artin stacks equipped with locally-free log structures.
@@ 3.4 Relative shifted log-symplectic forms
Recall that clasically a map $R\longrightarrow S$ of discrete commutative $\mathbb{k}$-algebras is finitely presented if $\mathsf{Hom}_{R/calg}(S,-)$ commutes with filtered colimits. Above we used the homotopical notion of being locally finitely presented, defined as follows:
@@
[33, Definition 2.16]  We say that a map $A\longrightarrow B$ in $\mathrm{cdga}^{\leq 0}$ is (homotopically) finitely presented if $\mathsf{Map}_{A/\mathrm{cdga}^{\leq 0}}(B,-)$ commutes with (homotopy) filtered colimits.
@@
A map $A\longrightarrow B$ in $\mathrm{cdga}^{\leq 0}$ is finitely presented ([35, 2.2.] and [22, Theorem 7.4.3.18]) if and only if
  1. 1.
@@
We will also need an appropriate notion of quasi-smoothness which we recall next:
@@
[29, Definition 2.12] The relative log cotangent complex of the log morphism of derived stacks $\pi:M\to S$ is the complex ${\mathbb{L}}^{log}_{M/S}:=\mathbb{L}_{\pi_{log}}$. Here $\mathbb{L}_{\pi_{log}}$ denotes the relative cotangent complex for the classifying morphism $\pi_{log}$.
@@
When $\pi:M\to S$ is a log quasi-smooth morphism from a derived Artin stack to a scheme $X$, then the quasi-smoothness of the map $\pi_{log}:M\longrightarrow\mathfrak{L}^{0}_{S,\mathsf{lf}}$ is equivalent to the fact that ${\mathbb{L}}^{log}_{M/S}$ is a perfect complex over $M$ of amplitude $[-1,1]$.
@@
$DR(M/N):=\operatorname{Sym}_{A}(\mathbb{L}_{M/N}[1]))=\oplus^{\infty}_{p=0}\wedge^{p}\mathbb{L}_{M/N}[p],$ (3.29)
@@
We also have the de Rham differential $d_{DR}:\wedge^{p}\mathbb{L}_{M/N}\longrightarrow\wedge^{p+1}\mathbb{L}_{M/N}$. The de Rham differential defines a mixed structure on $DR(M/N)$ with $\epsilon:=d_{DR}$. This makes $DR(M/N)$ into a mixed cdga. In addition to the internal cohomological grading $DR(M/N)$ has a second (form degree) “weight grading” defined as $DR(M/N)(p):=\wedge^{p}\mathbb{\mathbb{missing}}L_{M/N}[p]$. The weight grading and mixed structure on $DR(M/N)$ are also compatible with the multiplicative structure and make $DR(M/N)$ into a graded mixed cdga over $\mathbb{k}$. The degree and weight of the internal differential $d$ are $1$ and 00, respectively. The degree and weight of $\epsilon$ are $-1$ and $1$, respectively.
@@
Recall that for a mixed graded complex $F\in\epsilon-dg^{gr}$ of $\mathbb{k}$-vector spaces we have assoicated complexes $NC^{n}(F)(p):=\underset{i\geq 0}{\prod}F^{n-2i}(p+i)\in dg$ and a graded complex $NC^{w}(F):=\underset{p\in\mathbb{Z}}{\oplus}NC(F)(p)\in dg^{gr}$ of negative cyclic chains. Recall also that for any complex $E\in dg$ of $\mathbb{k}$-vector spaces we get a simplicial set $|E|=Map_{dg}(\mathbb{k},E)=$ the Dold-Kan functor applied to the $\tau_{\leq 0}$ truncation of $E$.
@@
The constructions $\boldsymbol{\mathcal{A}^{p}_{N}(-,n)}$ and $\boldsymbol{\mathcal{A}^{p,cl}_{N}(-,n)}$ give rise to $\infty$-functors $\mathbf{dAff}^{op}_{/N}\longrightarrow\mathbb{S}$ from the $\infty$-category $\mathbf{dAff}^{op}_{/N}$ of affine derived schemes over $N$ to the $\infty$-category $\mathbb{S}$ of simplicial sets.
@@
For $\operatorname{Spec}(A)\in\mathbf{dAff}_{/N}$, the simplicial set $\mathcal{A}^{p}_{N}(\operatorname{Spec}(A),n)$ (respectively $\mathcal{A}^{p,cl}_{N}(\operatorname{Spec}(A),n)$) is called the space of $p$-forms of degree $n$ on the derived stack $\operatorname{Spec}(A)$, relative to $N$ (respectively the space of closed $p$-forms of degree $n$ on the derived stack $\operatorname{Spec}(A)$, relative to $N$).
The two $\infty$-functors $\boldsymbol{\mathcal{A}^{p}_{N}(-,n)}$ and $\boldsymbol{\mathcal{A}^{p,cl}_{N}(-,n)}$ are derived prestacks on $\mathbf{dAff}_{/N}$. By [32, Proposition 1.11], we know that $\boldsymbol{\mathcal{A}^{p}_{N}(-,n)}$ and $\boldsymbol{\mathcal{A}^{p,cl}_{N}(-,n)}$ are derived stacks for the big étale site over $N$.
@@
    $\mathcal{A}_{log}^{p,cl}(M/S,n):=\mathsf{Map}_{{}_{\textbf{dSt}_{S,log}}}\left(M,\boldsymbol{\mathcal{A}}^{p,cl}_{\mathfrak{L}^{0}_{S,\mathsf{lf}}}(-,n)\right).$
@@
    A $2$-form $\omega\in\mathcal{A}_{log}^{2}(M/S,n)$ is non-degenerate if the induced map in $\mathrm{D}_{qcoh}(M)$
    $\Theta_{\omega}:\mathbb{T}^{log}_{M/S}\longrightarrow\mathbb{L}^{log}_{M/S}[n]$ (3.30)
    is a quasi-isomorphism. We denote by $\mathcal{A}_{log}^{2}(M/S,n)^{nd}$ the union of all the connected components of $\mathcal{A}_{log}^{2}(M/S,n)$ which is consist of non-degenerate relative logarithmic $2$-forms of degree $n$ on $M$.
@@
    $Symp_{log}(M/S,n):=\mathcal{A}_{log}^{2}(M/S,n)^{nd}\times^{h}_{\mathcal{A}_{log}^{2}(M/S,n)}\mathcal{A}_{log}^{2,cl}(M/S,n).$
4 The logarithmic Dolbeault moduli stack and its log symplectic structure
In this section, we will construct a logarithmic version of the relative Dolbeault moduli stack for the family of curves $\mathfrak{X}_{\mathfrak{M}}\rightarrow\mathfrak{M}$ defined at the end of section 3.2. We will show that the relative logarithmic Dolbeault moduli stack has a relative 00-shifted log-symplectic form over the spectrum of a discrete valuation ring $S$. Moreover, we will show that the relative log-symplectic form is an extension of Hitchin’s symplectic form on the generic fibre of the moduli stack over the spectrum of a discrete valuation ring. This was proved for moduli schemes in [9].
@@ 4 The logarithmic Dolbeault moduli stack and its log symplectic structure
$\omega_{C}=\ker\left[q_{*}\Omega_{\widetilde{C}}^{1}(\widetilde{D})\longrightarrow\underset{x\in D}{\bigoplus}\ \mathbb{k}_{x}\right],$ (4.1)
@@
    $s\mapsto\mathsf{Res}(s;x^{+})+\mathsf{Res}(s;x^{-}),$ (4.2)
@@
It is straightforward to check that it is a locally free sheaf of rank one. It is well known that this property persists in families. That is, for a family $f:\mathcal{X}\to\mathcal{S}$ of semi-stable/ pre-stable curves there is a well defined relative dualizing sheaf $\omega_{\mathcal{X}/\mathcal{S}}$ which is a line bundle on the total space of the family [39, 109.19 The relative dualizing sheaf]. In the case when $\mathcal{S}$ and $\mathcal{X}$ are both smooth, the relative dualizing sheaf is naturally identified with the sheaf of relative logarithmic forms along the fibers of $f$. Concretely, if $\boldsymbol{\mathfrak{d}}_{\mathcal{S}}\subset\mathcal{S}$ denotes the discriminant divisor of the map $f:\mathcal{X}\to\mathcal{S}$, and $\boldsymbol{\mathfrak{d}}_{\mathcal{X}}=f^{-1}\left(\boldsymbol{\mathfrak{d}}_{\mathcal{S}}\right)$ is the normal crossings divisor in $\mathcal{X}$ comprised of the singular fibers of $f$, then we have
@@
    $\mathcal{X}_{Dol/\mathcal{S}}:=\left[\widehat{\mathsf{T}_{\mathcal{X}/\mathcal{S}}}\times_{S}\mathcal{X}\rightrightarrows\mathcal{X}\right]$ (4.3)
@@
    $\mathcal{X}^{nil}_{Dol/\mathcal{S}}:=[\mathsf{T}_{\mathcal{X}/\mathcal{S}}\times_{\mathcal{S}}\mathcal{X}\rightrightarrows\mathcal{X}]$ (4.4)
@@
The proof follows verbatim [31, Proposition 5.1.2] after replacing the relative cotangent complex $\mathbb{L}_{\mathcal{X}/\mathcal{S}}$ with the relative dualizing sheaf $\omega_{\mathcal{X}/\mathcal{S}}$. ∎
@@
{Quasi-coherent $\left\{\begin{array}[]{@{}ll@{}}\text{Quasi-coherent sheaves}\\ \text{over}\leavevmode\nobreak\ B\widehat{\mathsf{T}}_{\mathcal{X}/\mathcal{S}}\end{array}\right\}\cong\left\{\begin{array}[]{@{}ll@{}}\text{Quasi-coherent sheaves over}\leavevmode\nobreak\ \mathcal{X}\leavevmode\nobreak\ \text{with}\\ \text{an action of the sheaf of algebras}\\ \operatorname{Sym}(\omega^{\vee}_{\mathcal{X}/\mathcal{S}})\end{array}\right\}$ (4.5)
Notice that $\operatorname{Sym}(\omega^{\vee}_{\mathcal{X}/\mathcal{S}})$ is a quadratic algebra. The quadratic dual algebra $(\operatorname{Sym}\leavevmode\nobreak\ \omega^{\vee}_{\mathcal{X}/\mathcal{S}})^{!}$ is isomorphic to the dg-algebra $\operatorname{Sym}(\omega_{\mathcal{X}/\mathcal{S}}[-1])$ with zero differential. Moreover, we have the following equivalence of categories ([30]).
{Quasi-coherent sheaves $\left\{\begin{array}[]{@{}ll@{}}\text{Quasi-coherent sheaves over}\leavevmode\nobreak\ \mathcal{X}\leavevmode\nobreak\ \text{with}\\ \text{an action of the sheaf of algebras}\\ \operatorname{Sym}(\omega^{\vee}_{\mathcal{X}/\mathcal{S}})\end{array}\right\}\cong\left\{\begin{array}[]{@{}ll@{}}\text{Quasi-coherent sheaves over}\\ \operatorname{Spec}_{\mathcal{X}}\leavevmode\nobreak\ \operatorname{Sym}\leavevmode\nobreak\ (\omega_{\mathcal{X}/\mathcal{S}}[-1])\end{array}\right\}$ (4.6)
@@
$B\widehat{\mathsf{T}}_{\mathcal{X}/\mathcal{S}}\cong\operatorname{Spec}_{\mathcal{X}}\leavevmode\nobreak\ \operatorname{Sym}(\omega_{\mathcal{X}/\mathcal{S}}[-1]).$ (4.7)
@@
{Vector $\left\{\begin{array}[]{@{}ll@{}}\text{Vector bundles}\\ \text{over}\leavevmode\nobreak\ B\widehat{\mathsf{T}}_{\mathcal{X}/\mathcal{S}}\end{array}\right\}\cong\left\{\begin{array}[]{@{}ll@{}}\text{Vector bundles over}\leavevmode\nobreak\ \mathcal{X}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{with an}\\ \text{ action of the sheaf of algebras}\\ ({\operatorname{Sym}(\omega_{\mathcal{X}/\mathcal{S}}}))^{\vee}\cong\widehat{\operatorname{Sym}}(\omega^{\vee}_{\mathcal{X}/\mathcal{S}})\end{array}\right\}\cong\left\{\begin{array}[]{@{}ll@{}}\text{nilpotent Higgs}\\ \text{bundles over}\leavevmode\nobreak\ \mathcal{X}\end{array}\right\}$ (4.8)
@@
$\mathcal{X}_{Dol/\mathcal{S}}$ is $\mathcal{O}$-compact and $\mathcal{O}$-oriented over $\mathcal{S}$. Hence, $\mathsf{Map}_{\mathcal{S}}(\mathcal{X}_{Dol/\mathcal{S}},BGL_{n}(\mathcal{O}_{\mathcal{S}}))$ has a 00-shifted relative symplectic structure over $\mathcal{S}$.
@@
${\mathcal{X}_{Dol/\mathcal{S}}}$${\mathcal{X}}$${\mathcal{S}}$$\scriptstyle{q}$$\scriptstyle{p}$$\scriptstyle{r}$ (4.9)
@@
$Rq_{*}\mathcal{O}_{\mathcal{X}_{Dol/\mathcal{S}}}\cong\mathcal{O}_{\mathcal{X}}\oplus\omega_{\mathcal{X}/\mathcal{S}}[-1],$ (4.10)
@@
$\eta\in(Rp_{*}\mathcal{O}_{\mathcal{X}_{Dol/\mathcal{S}}})^{\vee}[-2]=R\mathcal{H}om_{\mathcal{S}}(Rp_{*}\mathcal{O}_{\mathcal{X}_{Dol/\mathcal{S}}},\mathcal{O}_{\mathcal{S}}[-2]).$ (4.11)

### v3

@@ 3.3.1. Relative logarithmic cotangent complex
This implies that the image of $\sum^{n+1}_{i=1}\lambda_{i}\partial_{t_{i}}]\in T_{(1,\cdots,1)}(\mathbb{G}_{m})^{n}=\mathbb{k}^{n}$ (notice that $\sum^{n+1}_{i=1}\lambda_{i}=0$) is $\sum^{n+1}_{i=1}\lambda_{i}\cdot x_{i}\cdot\partial_{x_{i}}=\sum^{n+1}_{i=1}\lambda_{i}\cdot\partial_{\log\penalty 10000\ x_{i}}$. Therefore, we get a map
(3.24) $d(Or)_{p}:T_{(1,\cdots,1)}(\mathbb{G}_{m})^{n}=\mathbb{k}^{n}\longrightarrow(T\mathbb{A}^{n+1}(-\log\penalty 10000\ \partial A^{n+1}))_{p}=\mathbb{k}^{n+1}$
(3.25) $(\sum^{n+1}_{i=1}\lambda_{i}\cdot\partial_{t_{i}})\mapsto\sum^{n+1}_{i=1}\lambda_{i}\cdot\partial_{\log\penalty 10000\ x_{i}}$
The map above is clearly injective. Therefore, the dual of this map
(3.26) $(d(Or)_{p})^{\vee}:\Omega^{1}_{p}\mathbb{A}^{n+1}(\log\penalty 10000\ \partial A^{n+1})\longrightarrow\Omega^{1}_{(1,\cdots,1)}(\mathbb{G}_{m})^{n}$
@@
(3.27) $\Omega_{S}^{1}(\log o)\hookrightarrow\Omega_{S[n]}^{1}(\log\penalty 10000\ \penalty 10000\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}$
is 0, and that the morphism $\Omega_{S[n]/S}^{1}(\log\penalty 10000\ \penalty 10000\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}$ is an isomorphism.
@@
(3.28) $\big[\Omega_{S[n]/S}^{1}(\log\penalty 10000\ \penalty 10000\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}\big]$
is precisely the relative log cotangent complex of the morphism $\big[S[n]/G[n]\big]\longrightarrow S$ pulled back to $S[n]$. But since the morphism $\Omega_{S[n]/S}^{1}(\log\penalty 10000\ \penalty 10000\ \partial S[n])\longrightarrow\mathcal{O}_{S[n]}^{\oplus n}$ is an isomorphism, the relative log-cotangent complex of the morphism $\big[S[n]/G[n]\big]\longrightarrow S$ is equivalent to 0. ∎
3.4. Relative shifted log-symplectic forms
In this subsection, we will recall the definition of relative shifted symplectic forms for a locally of finite presentation morphism of derived Artin stacks. We will then define relative shifted log-symplectic forms for certain logarithmic morphisms of derived Artin stacks equipped with locally-free log structures.
@@ 3.4. Relative shifted log-symplectic forms
It is well-known that a map $R\longrightarrow S$ of commutative $\mathbb{k}$-algebras is finitely presented if and only if $\mathsf{Hom}_{R/calg}(S,-)$ commutes with filtered colimits. Above we used the homotopical notion of being locally finitely presented, defined as follows:
@@
[30, Definition 2.16] We say that a map $A\longrightarrow B$ in $\mathrm{cdga}^{\leq 0}$ is (homotopically) finitely presented if $\mathsf{Map}_{A/\mathrm{cdga}^{\leq 0}}(B,-)$ commutes with (homotopy) filtered colimits.
@@
A map $A\longrightarrow B$ in $\mathrm{cdga}^{\leq 0}$ is finitely presented ([30, Definition 2.16] and [23, Theorem 7.4.3.18]) if and only if
  1. (1)
@@
We will also need an appropriate notion of quasi-smoothness, which we recall next:
@@
[33, Definition 2.12] The relative log cotangent complex of the log morphism of derived stacks $\pi:M\to S$ is the complex ${\mathbb{L}}^{\log}_{M/S}:=\mathbb{L}_{\pi_{\log}}$. Here $\mathbb{L}_{\pi_{\log}}$ denotes the relative cotangent complex for the classifying morphism $\pi_{\log}$.
@@
When $\pi:M\to S$ is a log quasi-smooth morphism from a derived Artin stack to a scheme $X$, then the quasi-smoothness of the map $\pi_{\log}:M\longrightarrow\mathfrak{L}^{0}_{S,\mathsf{lf}}$ is equivalent to the fact that ${\mathbb{L}}^{\log}_{M/S}$ is a perfect complex over $M$ of amplitude [1,1][-1,1].
@@
(3.29) $DR(M/N):=\operatorname{Sym}_{A}(\mathbb{L}_{M/N}[1]))=\oplus^{\infty}_{p=0}\wedge^{p}\mathbb{L}_{M/N}[p],$
@@
We also have the de Rham differential $d_{DR}:\wedge^{p}\mathbb{L}_{M/N}\longrightarrow\wedge^{p+1}\mathbb{L}_{M/N}$. The de Rham differential defines a mixed structure on $DR(M/N)$ with $\epsilon:=d_{DR}$. This makes $DR(M/N)$ into a mixed cdga (we include the definition below). In addition to the internal cohomological grading $DR(M/N)$ has a second (form degree) “weight grading” defined as $DR(M/N)(p):=\wedge^{p}\mathbb{\mathbb{}}L_{M/N}[p]$. The weight grading and mixed structure on $DR(M/N)$ are also compatible with the multiplicative structure and make $DR(M/N)$ into a graded mixed cdga over $\mathbb{k}$. The degree and weight of the internal differential $d$ are $1$ and 0, respectively. The degree and weight of $\epsilon$ are 1-1 and $1$, respectively. For the definitions of mixed complexes, graded mixed complexes and mixed cdga’s we refer the reader to [29, Section 1.1].
@@
Recall that for a mixed graded complex $F\in\epsilon-dg^{gr}$ of $\mathbb{k}$-vector spaces we have assoicated complexes $NC^{n}(F)(p):=\underset{i\geq 0}{\prod}F^{n-2i}(p+i)\in dg$ and a graded complex $NC^{w}(F):=\underset{p\in\mathbb{Z}}{\oplus}NC(F)(p)\in dg^{gr}$ of negative cyclic chains. Recall also that for any complex $E\in dg$ of $\mathbb{k}$-vector spaces we get a simplicial set $|E|=\text{Map}_{dg}(\mathbb{k},E)=$ the Dold-Kan functor applied to the $\tau_{\leq 0}$ truncation of $E$.
@@
Again by precomposing with cofibrant replacements, the constructions $\mathcal{A}^{p}_{N}(-,n)$ and $\mathcal{A}^{p,cl}_{N}(-,n)$ give rise to $\infty$-functors $\mathbf{dAff}^{op}_{/N}\longrightarrow\mathbb{S}$ from the $\infty$-category $\mathbf{dAff}^{op}_{/N}$ of affine derived schemes over $N$ to the $\infty$-category $\mathbb{S}$ of simplicial sets. We denote these $\infty$-functors by $\boldsymbol{\mathcal{A}^{p}_{N}(-,n)}$ and $\boldsymbol{\mathcal{A}^{p,cl}_{N}(-,n)}$, respectively.
@@
For $\boldsymbol{\operatorname{Spec}}(A)\in\mathbf{dAff}_{/N}$, the simplicial set $\mathcal{A}^{p}_{N}(\boldsymbol{\operatorname{Spec}}(A),n)$ (respectively, $\mathcal{A}^{p,cl}_{N}(\boldsymbol{\operatorname{Spec}}(A),n)$) is called the space of $p$-forms of degree $n$ on the derived stack $\boldsymbol{\operatorname{Spec}}(A)$, relative to $N$ (respectively, the space of closed $p$-forms of degree $n$ on the derived stack $\boldsymbol{\operatorname{Spec}}(A)$, relative to $N$).
The two $\infty$-functors $\boldsymbol{\mathcal{A}^{p}_{N}(-,n)}$ and $\boldsymbol{\mathcal{A}^{p,cl}_{N}(-,n)}$ are derived prestacks on $\mathbf{dAff}_{/N}$. By [29, Proposition 1.11], we know that $\boldsymbol{\mathcal{A}^{p}_{N}(-,n)}$ and $\boldsymbol{\mathcal{A}^{p,cl}_{N}(-,n)}$ are derived stacks for the big étale site over $N$.
@@
    $\mathcal{A}_{\log}^{p,cl}(M/S,n):=\mathsf{Map}_{{}_{\textbf{dSt}_{S,log}}}\left(M,\boldsymbol{\mathcal{A}}^{p,cl}_{\mathfrak{L}^{0}_{S,\mathsf{lf}}}(-,n)\right).$
@@
    A $2$-form $\omega\in\mathcal{A}_{\log}^{2}(M/S,n)$ is non-degenerate if the induced map in $\mathrm{D}_{qcoh}(M)$
    (3.30) $\Theta_{\omega}:\mathbb{T}^{\log}_{M/S}\longrightarrow\mathbb{L}^{\log}_{M/S}[n]$
    is a quasi-isomorphism. We denote by $\mathcal{A}_{\log}^{2}(M/S,n)^{nd}$ the union of all the connected components of $\mathcal{A}_{\log}^{2}(M/S,n)$ which is consist of non-degenerate relative logarithmic $2$-forms of degree $n$ on $M$. Here, $\mathbb{T}^{\log}_{M/S}$ denotes the dual of $\mathbb{L}^{\log}_{M/S}$, and we call it the relative logarithmic tangent complex.
@@
    $Symp_{\log}(M/S,n):=\mathcal{A}_{\log}^{2}(M/S,n)^{nd}\times^{h}_{\mathcal{A}_{\log}^{2}(M/S,n)}\mathcal{A}_{\log}^{2,cl}(M/S,n).$
4. The logarithmic Dolbeault moduli stack and its log symplectic structure
In this section, we will construct a logarithmic version of the relative Dolbeault moduli stack for the family of curves $\mathfrak{X}_{\mathfrak{M}}\rightarrow\mathfrak{M}$ defined at the end of section 3.2. We will show that the relative logarithmic Dolbeault moduli stack has a relative 0-shifted log-symplectic form over the spectrum of a discrete valuation ring $S$. Moreover, we will show that the relative log-symplectic form is an extension of Hitchin’s symplectic form on the generic fibre of the moduli stack over the spectrum of a discrete valuation ring. This was proved for moduli schemes in [10].
@@ 4. The logarithmic Dolbeault moduli stack and its log symplectic structure
(4.1) $\omega_{C}=\ker\left[q_{*}\Omega_{\widetilde{C}}^{1}(\widetilde{D})\longrightarrow\underset{x\in D}{\bigoplus}\ \mathbb{k}_{x}\right],$
@@
    (4.2) $s\mapsto\mathsf{Res}(s;x^{+})+\mathsf{Res}(s;x^{-}),$
@@
It is straightforward to check that it is a locally free sheaf of rank one. It is well known that this property persists in families. That is, for a family $f:\mathcal{X}\to\mathcal{S}$ of semi-stable/ pre-stable curves there is a well defined relative dualizing sheaf $\omega_{\mathcal{X}/\mathcal{S}}$ which is a line bundle on the total space of the family [41, 109.19 The relative dualizing sheaf]. In the case when $\mathcal{S}$ and $\mathcal{X}$ are both smooth, the relative dualizing sheaf is naturally identified with the sheaf of relative logarithmic forms along the fibers of $f$. Concretely, if $\boldsymbol{\mathfrak{d}}_{\mathcal{S}}\subset\mathcal{S}$ denotes the discriminant divisor of the map $f:\mathcal{X}\to\mathcal{S}$, and $\boldsymbol{\mathfrak{d}}_{\mathcal{X}}=f^{-1}\left(\boldsymbol{\mathfrak{d}}_{\mathcal{S}}\right)$ is the normal crossings divisor in $\mathcal{X}$ comprised of the singular fibers of $f$, then we have
@@
    (4.3) $\mathcal{X}_{Dol/\mathcal{S}}:=\left[\widehat{\mathsf{T}_{\mathcal{X}/\mathcal{S}}}\times_{S}\mathcal{X}\rightrightarrows\mathcal{X}\right]$
@@
    (4.4) $\mathcal{X}^{nil}_{Dol/\mathcal{S}}:=[\mathsf{T}_{\mathcal{X}/\mathcal{S}}\times_{\mathcal{S}}\mathcal{X}\rightrightarrows\mathcal{X}]$
@@
The proof follows verbatim [32, Proposition 5.1.2] after replacing the relative cotangent complex $\mathbb{L}_{\mathcal{X}/\mathcal{S}}$ with the relative dualizing sheaf $\omega_{\mathcal{X}/\mathcal{S}}$. ∎
@@
[…]
Notice that $\operatorname{Sym}(\omega^{\vee}_{\mathcal{X}/\mathcal{S}})$ is a quadratic algebra. The quadratic dual algebra $(\operatorname{Sym}\penalty 10000\ \omega^{\vee}_{\mathcal{X}/\mathcal{S}})^{!}$ is isomorphic to the dg-algebra $\operatorname{Sym}(\omega_{\mathcal{X}/\mathcal{S}}[-1])$ with zero differential. Moreover, we have the following equivalence of categories ([31]).
[…]
@@
(4.7) $B\widehat{\mathsf{T}}_{\mathcal{X}/\mathcal{S}}\cong\operatorname{Spec}_{\mathcal{X}}\penalty 10000\ \operatorname{Sym}(\omega_{\mathcal{X}/\mathcal{S}}[-1]).$
@@
[…]
@@
$\mathcal{X}_{Dol/\mathcal{S}}$ is $\mathcal{O}$-compact and $\mathcal{O}$-oriented over $\mathcal{S}$. Hence, $\mathsf{Map}_{\mathcal{S}}(\mathcal{X}_{Dol/\mathcal{S}},BGL_{n}(\mathcal{O}_{\mathcal{S}}))$ has a 0-shifted relative symplectic structure over $\mathcal{S}$.
@@
(4.9) $\mathcal{X}_{Dol/\mathcal{S}}$$\mathcal{X}$$\mathcal{S}$$\scriptstyle{q}$$\scriptstyle{p}$$\scriptstyle{r}$
@@
(4.10) $Rq_{*}\mathcal{O}_{\mathcal{X}_{Dol/\mathcal{S}}}\cong\mathcal{O}_{\mathcal{X}}\oplus\omega_{\mathcal{X}/\mathcal{S}}[-1],$
@@
(4.11) $\eta\in(Rp_{*}\mathcal{O}_{\mathcal{X}_{Dol/\mathcal{S}}})^{\vee}[-2]=R\mathcal{H}om_{\mathcal{S}}(Rp_{*}\mathcal{O}_{\mathcal{X}_{Dol/\mathcal{S}}},\mathcal{O}_{\mathcal{S}}[-2]).$